### v1

# Quantum entanglement, coherence and Dark Matter

Abstract
In this paper we consider the effect of quantum entanglement and coherence on the radiated intensity from a gas and its absorption capacity at thermal equilibrium or, more generally, under conditions where no population inversion exists. As was shown by Dicke (1954), although entangled states and coherence can lead to superradiance for specific modes of radiation, they can also bring subradiance through significant energy trapping in slow and dark states. While a finite separation between the atoms composing the gas will cause leaking of the trapped energy, we show how the combination of thermal equilibrium and quantum coherence mitigates this effect and leads to significantly reduced radiation intensity from the gas, rendering it dark and collision-less. Furthermore, we show how under the same conditions absorption of a radiation field incident on the gas can lead to higher attenuation levels than those predicted with Beer’s law. Beer’s law is recovered in the limit of complete decoherence. We apply our analysis to the atomic hydrogen 21 cm line and, considering the gas densities expected in Dark Matter halos, we find that quantum entanglement and coherence can potentially account for some of the Dark Matter known exist in these environments.
I Introduction
It has been known since Dicke’s original treatment of the spontaneous emission of coherent radiation [1] that the interaction between atoms composing a gas  [^1] Although we focus on atoms in this paper, our analysis and discussions apply to molecular gases as well. through their common electromagnetic field leads to the formation of entangled quantum mechanical states, with significant implications for the spontaneous emission rates of photons emanating from the gas. As Dicke pointed out, atoms should not be considered independently; rather, the entire gas must be treated as a single quantum mechanical system. While the material covered in his paper was broad, touching on several phenomena involving coherence and entanglement, much of the early work it inspired focused on superradiance.
Superradiance is a fundamental radiation process [3] in which the cooperative behavior of the atoms composing the gas leads to powerful bursts of coherent radiation. Indeed, superradiance can be thought of as “collective spontaneous emission,” where, instead of spontaneously emitting photons individually, the atoms cooperate in a coherent cascade of photon emission. Although the first experimental verification of superradiance occurred almost 20 years after Dicke’s work [4], astronomers remained largely unaware of this phenomenon and the intense research being conducted in the quantum optics community. This remained the case despite the earlier discovery of astronomical masers in the interstellar medium [5] and the complementarity between the maser action and superradiance [3, 6]. It is only recently that superradiance has been studied in an astronomical context and used to explain intense flares from maser-hosting regions in diverse environments [7, 8, 9, 10, 11, 12, 6, 13, 14, 15]. Beyond improving our understanding and modeling of these systems, these studies have importantly established the presence of the type of coherence proposed by Dicke in astronomical media.
Dicke’s 1954 paper [1] also explicitly discussed another coherent radiation process, which later became known as subradiance. In contrast to superradiance, subradiance describes the suppression of photon emission compared to the typical spontaneous emission process. Dicke analyzed the trapping of energy in slow and dark states and its consequences for the radiation intensity emanating from a gas. Subradiance was experimentally verified later than superradiance [16] and has only recently been observed in large clouds of cold atoms [17]. The latter experimental result is particularly relevant to us as we explore the possible existence of subradiant systems in astrophysical contexts.
More specifically, we ask the following question: Given that superradiance is known to occur in diverse astronomical environments, should we not also expect subradiance to be realized? We aim to answer this by exploring a possible link between subradiance and the presence of Dark Matter in the universe – one of the deepest mysteries in astrophysics since its manifestation in galactic rotation curves was discovered [18].
Our paper is structured as follows. In Sec. II, we review key material from Dicke’s 1954 paper [1], focusing on energy trapping. In Sec. III, we extend this analysis by introducing equilibrium conditions within the framework of the master equation. In Sec. IV, we apply the material developed in the previous sections to astrophysical contexts using the Maxwell-Bloch equations. Specifically, we focus on the 21cm line of atomic hydrogen to investigate the potential realization of subradiance in environments with densities consistent with those of galactic Dark Matter halos. Finally, we discuss our results in Sec. V, while detailed calculations are provided in three appendices at the end.
II Entangled quantum states and coherent behavior
In this section we revisit parts of Dicke’s work and focus on the trapping of energy within an atomic ensemble brought about by the existence of Dicke states of differing symmetries. Accordingly, we follow Dicke [1] and express the Hamiltonian for the gas with
$\hat{H}=\hat{H}_{0}+\hbar\omega_{0}\sum_{j}\hat{R}^{3}_{j}+\hat{V},$ (1)
where $\hat{H}_{0}$ accounts for the translational motions of the atoms’ center-of-mass and their interactions, the second term for their internal energy and $\hat{V}$ for dipole interactions with the electromagnetic field. For simplicity we approximate the atoms as two-level systems at resonance with each other (i.e., moving at the same velocity) with an internal energy difference $\hbar\omega_{0}$; the summation in equation (1) is on all atoms. That is, if |$\ket{a}$ and |$\ket{b}$ are the lower and upper atomic states, respectively, then for the $j^{\mathrm{th}}$ atom $\hat{R}^{3}_{j}=\frac{1}{2}\left(\left|b\middle\rangle\middle\langle b\right|-\left|a\middle\rangle\middle\langle a\right|\right)$.
For our problem the interaction term reduces to
$\hat{V}=-i\sum_{k}\frac{1}{2}\hbar\Omega_{k}\left(\hat{R}^{+}_{k}\hat{a}_{k}-\hat{R}^{-}_{k}\hat{a}^{\dagger}_{k}\right),$ (2)
where the summation on $k$ is on the radiation modes, and $\hat{a}_{k}$ and $\hat{a}^{\dagger}_{k}$ are the photon annihilation and creation operators. In equation (2) the Rabi frequency is given by
$\Omega_{k}=\frac{2\mu\mathcal{E}_{0}}{\hbar c}\left(\bm{\epsilon}_{\mu}\cdot\bm{\epsilon}_{k}^{\perp}\right)$ (3)
for a magnetic dipole transition, where $\mu$ and $\bm{\epsilon}_{\mu}$ are the magnitude of the magnetic dipole moment and its associated unit vector (assumed real), while that for the radiation mode’s polarization is (also real and) denoted by $\bm{\epsilon}_{k}^{\perp}$ (i.e., $\bm{k}=k\bm{\epsilon}_{k}$ and $\bm{\epsilon}_{k}\cdot\bm{\epsilon}_{k}^{\perp}=0$). The one-photon electric field is
$\mathcal{E}_{0}=\sqrt{\frac{\hbar\omega_{0}}{2\epsilon_{0}\mathcal{V}}}$ (4)
with $\epsilon_{0}$ the permittivity of vacuum and $\mathcal{V}$ the volume of quantization. We focus on magnetic dipole transitions in view of our application to the atomic hydrogen 21 cm line in Sec. IV. For an electrical dipole transition one has to substitute $\mu\rightarrow dc$ and $\bm{\epsilon}_{\mu}\rightarrow\bm{\epsilon}_{d}$ in equation (3), with $d$ the magnitude of the electric dipole moment.
The phase-matched raising and lowering operators introduced in equation (2) are defined by
$\hat{R}^{\pm}_{k}=\sum_{j}\hat{R}^{\pm}_{j}e^{\pm i\bm{k}\cdot\mathbf{r}_{j}}$ (5)
with $\mathbf{r}_{j}$ the position of atom $j$. For the single-atom raising and lowering operators we have $\hat{R}^{+}_{j}=\left|b\middle\rangle\middle\langle a\right|$ and $\hat{R}^{-}_{j}=\left|a\middle\rangle\middle\langle b\right|$ while, as we assume resonance, we write $k=\left|\mathbf{k}\right|=\omega_{0}/c$.
II.1 The two-atom problem
Let us consider the problem of two identical two-level atoms located a distance $z_{0}$ from one another; Atom 1 at $-z_{0}/2$ and Atom 2 at $z_{0}/2$ on the $z$-axis. For this system
$\hat{R}^{\pm}_{k}=\hat{R}^{\pm}_{1}e^{\mp i\frac{1}{2}kz_{0}\cos\theta}+\hat{R}^{\pm}_{2}e^{\pm i\frac{1}{2}kz_{0}\cos\theta}$ (6)
with $\theta$ the angle of k relative to the $z$-axis.
Given that single-atom transitions take place between the lower |$\Ket{a}$ and upper |$\Ket{b}$ levels of energies $-\hbar\omega_{0}/2$ and $\hbar\omega_{0}/2$, respectively, appropriate internal states for the two-atom problem consist of [1, 19] (see Appendix C)
|$\Ket{1,1}=\Ket{bb}$ (7)
|$\Ket{1,0}_{\theta}=\frac{1}{\sqrt{2}}\left(e^{i\frac{1}{2}kz_{0}\cos\theta}\Ket{ab}+e^{-i\frac{1}{2}kz_{0}\cos\theta}\Ket{ba}\right)$ (8)
|$\Ket{1,-1}=\Ket{aa}$ (9)
|$\Ket{0,0}_{\theta}=\frac{1}{\sqrt{2}}\left(e^{i\frac{1}{2}kz_{0}\cos\theta}\Ket{ab}-e^{-i\frac{1}{2}kz_{0}\cos\theta}\Ket{ba}\right).$ (10)
Following Dicke [1], the kets |$\Ket{r,m}$ are defined with the ‘cooperative’ number $r$ and ‘inversion’ number $m$, which are analogous to spin-1/2 quantum numbers, with $r\geq 0$ and $\left|m\right|\leq r$ (note that |$\Ket{r,m}_{\theta}=\Ket{r,m}$ when $\cos\theta=0$).
Given equations (7)-(10) it is straightforward to calculate radiative transition rates between the different states upon applying Fermi’s Golden Rule with equations (2), (3) and (6). When $kz_{0}\neq 0$ the selection rules are then found to be $\Delta r=0,\pm 1$ and $\Delta m=\pm 1$, and the spontaneous emission rates per unit solid angle
$\frac{d\gamma_{1,1\rightarrow 1,0_{\theta}}}{d\Omega}=2\frac{d\Gamma}{d\Omega}$ (11)
$\frac{d\gamma_{1,1\rightarrow 0,0_{\theta}}}{d\Omega}=0$ (12)
$\frac{d\gamma_{1,0_{\theta}\rightarrow 1,-1}}{d\Omega^{\prime}}=2\cos^{2}\left[\frac{1}{2}kz_{0}\left(\cos\theta^{\prime}-\cos\theta\right)\right]\frac{d\Gamma}{d\Omega^{\prime}}$ (13)
$\frac{d\gamma_{0,0_{\theta}\rightarrow 1,-1}}{d\Omega^{\prime}}=2\sin^{2}\left[\frac{1}{2}kz_{0}\left(\cos\theta^{\prime}-\cos\theta\right)\right]\frac{d\Gamma}{d\Omega^{\prime}}$ (14)
with $d\Gamma/d\Omega$ the spontaneous emission rate per unit solid angle of a single independent atom, and $\theta$ and $\theta^{\prime}$ denoting the orientations of the first and second emitted photons, respectively, relative to the $z$-axis.
The signatures of entanglement and coherence are displayed in equations (13)-(14). That is, the enhancement of the transitions rates by a factor of ‘2’ is due to the entangled nature of the states given in equations (8) and (10), while the disappearance of the argument in the cosine and sine functions stems from coherence between the two emitted photons (i.e., when $\cos\theta^{\prime}=\cos\theta$). We therefore find the existence of an angular correlation between successively emitted photons. More precisely, if the system starts in the |$\Ket{1,1}$ state, then after the emission of the first photon it will for sure end up in the symmetric |$\Ket{1,0}_{\theta}$ intermediate state. From there, we see that the second photon has a higher probability of being emitted in a direction where $\cos\theta^{\prime}=\cos\theta$ to reach the |$\Ket{1,-1}$ ground state [19].
It is customary in studies of superradiance to set the initial conditions such that the ensemble of atoms is initially in a state of complete inversion (i.e., all atoms are in the excited state |$\Ket{b}$, as done in the previous paragraph [20]). However, in later sections we will be interested in situations where the system is in thermal equilibrium where all the states for a given $m$ value are equally populated. We will thus, from now on, focus on the transition rates between the $m=0$ and $m=-1$ states given in equations (13)-(14). To simplify our discussion we set $\cos\theta=0$ and integrate these transition rates over $d\Omega^{\prime}$ to find
$\gamma_{1,0_{\theta}\rightarrow 1,-1}=\Gamma\left[1+F\left(kz_{0}\right)\right]$ (15)
$\gamma_{0,0_{\theta}\rightarrow 1,-1}=\Gamma\left[1-F\left(kz_{0}\right)\right]$ (16)
with
[…]
$\left.+\left[1-3\left(\bm{\epsilon}_{\mu}\cdot\bm{\epsilon}_{z}\right)^{2}\right]\left[\frac{\cos\left(kz_{0}\right)}{\left(kz_{0}\right)^{2}}-\frac{\sin\left(kz_{0}\right)}{\left(kz_{0}\right)^{3}}\right]\right\}$ (17)
and
$\Gamma=\frac{\mu_{0}\omega_{0}^{3}\mu^{2}}{3\pi\hbar c^{3}}$ (18)
the single-atom free-space spontaneous emission rate with $\mu_{0}$ the permeability of vacuum (for an electric dipole transition we again substitute $\mu\rightarrow dc$).
We consider initial conditions where the states |$\Ket{1,0}_{\theta}$ and |$\Ket{0,0}_{\theta}$ have equal probabilities of occupation (of $1/2$) but with uncorrelated probability amplitude coefficients. Under these conditions, the radiation intensity can be evaluated with
$I\left(t\right)=\hbar\omega_{0}\left[P_{1,0_{\theta}}\left(t\right)\gamma_{1,0_{\theta}\rightarrow 1,-1}+P_{0,0_{\theta}}\left(t\right)\gamma_{0,0_{\theta}\rightarrow 1,-1}\right],$ (19)
where $P_{r,m_{\theta}}\left(t\right)$ is the probability of being in state |$\Ket{r,m}_{\theta}$ at time $t$; the different $P_{r,m_{\theta}}\left(t\right)$ are readily calculated from equations (15)-(16) and the initial probabilities. We thus have
$I\left(t\right)=\frac{1}{2}\hbar\omega_{0}\Gamma e^{-\Gamma t}$$\left\{\left[1+F\left(kz_{0}\right)\right]e^{-\Gamma F\left(kz_{0}\right)t}\right.$
$\left.+\left[1-F\left(kz_{0}\right)\right]e^{\Gamma F\left(kz_{0}\right)t}\right\}.$ (20)
It is instructive to look at the characteristics of the system at opposite limits of separation between the two atoms. More precisely, in the limit $kz_{0}\rightarrow\infty$ where $F\left(kz_{0}\right)\rightarrow 0$ we find
$I_{\infty}\left(t\right)=\hbar\omega_{0}\Gamma e^{-\Gamma t},$ (21)
which is exactly the intensity from a single independent atom. In other words, when separated by an infinite distance the two entangled atoms will spontaneously emit one photon with a probability of unity over a time-scale $\Gamma^{-1}$ (when only one is in the excited state), just as an independent excited atom would.
On the other hand, in the idealized small sample limit $kz_{0}=0$ when $F\left(kz_{0}\right)=1$ the intensity becomes
$I_{\mathrm{ss}}\left(t\right)=\frac{1}{2}\hbar\omega_{0}\left(2\Gamma\right)e^{-2\Gamma t}.$ (22)
The two-atom system will, on average, emit a photon half of the time but at twice the single-atom rate.
Another way to look at this is by considering the energy trapping efficiency defined as
[…]
$=0.5.$ (23)
Again, we find that half of the internal energy initially stored in the system remains within it in the steady-state.
The small sample behavior is readily understood from equations (13)-(14)(or 15-17) since in this case coherence happens for all radiation modes, resulting in the doubling and cancellation of corresponding transition rates. The latter implies that |$\Ket{0,0}_{\theta}$ is a dark state and that its initial occupation probability is preserved at all times. This is where energy trapping originates. On the other hand, the behavior for $kz_{0}\rightarrow\infty$ is a global characteristic in the sense that it can only be discerned in the solid-angle-integrated transition rate of equation (16) through the function $F\left(kz_{0}\right)$. Even at very large separations some radiation modes will be coherent, but when considering all modes the two atoms behave as if they were independent.
This last observation is important since external factors can sometimes render specific radiation modes dominant over the rest. For example, elongated (e.g., cylindrical) geometries are often considered in studies of superradiance since in such cases radiation along the symmetry axis is favored [19, 21, 20, 6]. Similarly, a gas in the interstellar medium initially hosting a population inversion will be more likely to radiate in a direction where velocity coherence is accentuated, leading to a transient superradiance response [6, 14]. Such considerations will be central to the material presented in Sec. IV.
Finally, we note that the converse of superradiance takes place when the two-atom system is initially in the ground |$\Ket{1,-1}$ state and subjected to an incident one-photon field. In this case, superabsorption happens and the transition rates of equations (11)-(14) are reversed [22]. That is,
$\frac{d\gamma_{1,-1\rightarrow 1,0_{\theta}}}{d\Omega}=2\frac{d\Gamma}{d\Omega}$ (24)
$\frac{d\gamma_{1,-1\rightarrow 0,0_{\theta}}}{d\Omega}=0$ (25)
$\frac{d\gamma_{1,0_{\theta}\rightarrow 1,1}}{d\Omega^{\prime}}=2\cos^{2}\left[\frac{1}{2}kz_{0}\left(\cos\theta^{\prime}-\cos\theta\right)\right]\frac{d\Gamma}{d\Omega^{\prime}}$ (26)
$\frac{d\gamma_{0,0_{\theta}\rightarrow 1,1}}{d\Omega^{\prime}}=2\sin^{2}\left[\frac{1}{2}kz_{0}\left(\cos\theta^{\prime}-\cos\theta\right)\right]\frac{d\Gamma}{d\Omega^{\prime}}$ (27)
with $\theta$ and $\theta^{\prime}$ denoting the orientations of the first and second absorbed photons, respectively, relative to the $z$-axis (the absorption of the second photon is also assumed to stem from a one-photon field). Here again, there is angular correlation, but this time between successively absorbed photons. Superabsorption will also play an important role when our model is applied to more general considerations in Sec. IV.
II.2 Generalization to $n$ atoms
For the two-atom problem only two permutation symmetries are available for the Dicke states: the totally symmetric species is shared by the triplet $r=1$ (equations 7-9), while the lone $r=0$ singlet is anti-symmetric (equation 10). In the more general case where the gas is composed of an arbitrary number of atoms $n$ a larger number of symmetries, tied to the cooperative number $r$, exist for the Dicke states. For example, it can be shown that a four-atom system would allow for five non-degenerate totally symmetric $r=2$ states, three levels of triply-degenerate $r=1$ states, and one level of doubly-degenerate $r=0$ dark states.
The general case for an arbitrary number of atoms $n$ is shown in Figure 1 along with the expected transition rates in the small sample (i.e., $kz_{0}\rightarrow 0$) or coherent limits (taken from Dicke’s original paper [1]). As can be seen in the figure, it is generally observed that, for a given $m$ value, transitions are allowed only between states sharing the same cooperation number $r$, while transition rates systematically decrease as one moves rightward to lower $r$ values. This behavior is consistent with that observed for the two-atom problem treated above.
Figure 1: The coherent radiation rates between Dicke states in the small sample (i.e., when $kr\rightarrow 0$) or coherent limits. The number of atom contained in the radiating gas is $n$, while the states within the different coherent cascades are identified through the cooperative number $20\leq r\leq n/$ 2 and the inversion number $-r\leq m\leq r$. The radiative intensity (or rate) of a single atom is denoted by $I_{0}$. Taken from Dicke’s original 1954 paper [1].
Although it is in principle possible to extend the spontaneous emission rates analysis for finite separations previously performed for the two-atom case (in Sec. II.1) to the four- or $n$-atom problems using appropriate Dicke states, the calculations quickly become prohibitive even for a restricted number of radiators. Such analysis would, for example, allow us to estimate the amount of energy trapped in the $n$-atom gas when starting with some initial conditions (e.g., equally populating the $m=0$ states) and establish any potential tendency for the radiation intensity as $n$ increases for a given spatial distribution.
It is, however, possible to determine the ratio of numbers of photons emitted in the small and infinite-size samples when $m=0$, as was done in equation (23) for the two-atom problem. For this, we use the known level of degeneracy of states as a function of the cooperative number $r$ [1]
$g_{r}=\frac{n!\left(2r+1\right)}{\left(n/2+r+1\right)!\left(n/2-r\right)!},$ (28)
with the fact that, for $n$ even, the number of photons emitted from $m=0$ for the small and infinite-size samples are $r$ and $n/2$, respectively. We thus find for the energy trapping efficiency 222for $n$ odd, we have for the energy trapping efficiency $\eta=1-\frac{2}{n-1}\left[\sum_{r=1/2}^{n/2}g_{r}\left(r-1/2\right)\right]/\left(\sum_{r=1/2}^{n/2}g_{r}\right)$ from $m=-1/2$.
$\eta=1-\frac{2}{n}\frac{\sum_{r=0}^{n/2}g_{r}r}{\sum_{r=0}^{n/2}g_{r}}.$ (29)
One can readily verify that $\eta=1/2$ for the two-atom case, while it is $0.69$ for $n=10$, $0.88$ for $n=100$, etc. It is therefore clear that the combination of coherence and the entanglement of a large number of atoms can potentially lead to a significant amount of energy trapping within the gas. We also note that this effect gets more pronounced as the system starts at lower $m$ values. For example, we find $\eta=1-1/n$ for equal initial occupation probabilities of the $m=-n/2+1$ states [1].
But one of our goals is to assess the level of energy trapping within the gas for finite and more realistic atomic separations in the interstellar medium, i.e., beyond the small sample limit as we have done so far. It is already obvious through our study of the two-atom system that the general case where $kz_{0}\neq 0$ will lead to a non-zero transition rate from the |$\Ket{0,0}_{\theta}$ ‘dark’ state (see equations 10 and 16). This has for implication that all the energy initially stored in the system will eventually leak away. That is, $\eta=0$ in the steady-state and no energy trapping is realized in the long run.
III Equilibrium conditions
To further our study of trapping and leaking of internal energy we move to density matrix framework. More precisely, we will make use of the master equation often used within the context of superradiance [20, 24]
$\frac{d\hat{\rho}}{dt}$ $=\frac{1}{i\hbar}\hbar\omega_{0}\sum_{j}\left[\hat{R}^{3}_{j},\hat{\rho}\right]-\frac{1}{i\hbar}\sum_{i\neq j}\Omega_{ij}\left(k\Delta r_{ij}\right)\left[\hat{R}^{+}_{i}\hat{R}^{-}_{j},\hat{\rho}\right]$
$-\frac{\Gamma}{2}\sum_{ij}F_{ij}\left(k\Delta r_{ij}\right)\left(\hat{R}^{+}_{i}\hat{R}^{-}_{j}\hat{\rho}+\hat{\rho}\hat{R}^{+}_{i}\hat{R}^{-}_{j}-2\hat{R}^{-}_{j}\hat{\rho}\hat{R}^{+}_{i}\right)$ (30)
with
$F_{ij}$ $\left(k\Delta r_{ij}\right)=\frac{3}{2}\left\{\left[1-\left(\bm{\epsilon}_{\mu}\cdot\bm{\epsilon}_{r}\right)^{2}\right]\frac{\sin\left(k\Delta r_{ij}\right)}{k\Delta r_{ij}}\right.$
$\left.+\left[1-3\left(\bm{\epsilon}_{\mu}\cdot\bm{\epsilon}_{r}\right)^{2}\right]\left[\frac{\cos\left(k\Delta r_{ij}\right)}{\left(k\Delta r_{ij}\right)^{2}}-\frac{\sin\left(k\Delta r_{ij}\right)}{\left(k\Delta r_{ij}\right)^{3}}\right]\right\}$ (31)
$\Omega_{ij}$ $\left(k\Delta r_{ij}\right)=\frac{3}{4}\left\{\left[1-\left(\bm{\epsilon}_{\mu}\cdot\bm{\epsilon}_{r}\right)^{2}\right]\frac{\cos\left(k\Delta r_{ij}\right)}{k\Delta r_{ij}}\right.$
$\left.-\left[1-3\left(\bm{\epsilon}_{\mu}\cdot\bm{\epsilon}_{r}\right)^{2}\right]\left[\frac{\sin\left(k\Delta r_{ij}\right)}{\left(k\Delta r_{ij}\right)^{2}}+\frac{\cos\left(k\Delta r_{ij}\right)}{\left(k\Delta r_{ij}\right)^{3}}\right]\right\}$ (32)
and where $\hat{\rho}$ is the density matrix for the atomic system, $\Delta r_{ij}=\left|\mathbf{r}_{i}-\mathbf{r}_{j}\right|$ is the distance between Atoms $i$ and $j$, $\bm{\epsilon}_{r}=\left(\mathbf{r}_{i}-\mathbf{r}_{j}\right)/\Delta r_{ij}$ and $\Gamma$ is given by equation (18).
For our two-atom problem the elements of the density matrix will be defined using the basis {|$\left\{\Ket{1}\equiv\Ket{1,0}\right.$, |$\Ket{2}\equiv\Ket{0,0}$, |$\left.\Ket{3}\equiv\Ket{1,-1}\right\}$, where |$\Ket{r,m}$ is given in equations (7)-(9) with $\cos\theta=0$. That is, we have $\rho_{ij}=\braket{i}{\hat{\rho}}{j}$. The solution of the master equation for the two-atom problem discussed in Sec. II.1 is presented in Figure 2 for $\Delta r=0.02\lambda$ with the black curves tracing the time evolution of the corresponding populations ($\lambda$ is the radiation wavelength). The initial conditions were set to $\rho_{11}\left(0\right)=\rho_{22}\left(0\right)=1/2$ and all other populations and coherences to zero, as well as $\bm{\epsilon}_{\mu}\cdot\bm{\epsilon}_{r}=0$.
As can be inferred from the figure, no energy is trapped in the system as the populations corresponding to the |$\Ket{1,0}$ and |$\Ket{0,0}$ states go to zero in the steady-state. The first (fast) state does so quickly (on a time-scale of $\Gamma t\sim 1/2$) while the second (dark) state decays on a much longer time-scale, but will still eventually go down to zero.
Figure 2: Solution of the two-atom system under the initial conditions $\rho_{11}\left(0\right)=\rho_{22}\left(0\right)=1/2$ and all other populations and coherences set to zero. The separation between the atoms was set to $\Delta r=0.02\lambda$. The time evolution of the relevant populations are traced using the legend for the corresponding states. Black curves: no equilibrium conditions are enforced and the populations corresponding to $m=0$ will eventually all relax to zero. No energy trapping takes place in the steady-state (i.e., when $\Gamma t\rightarrow\infty$). Colored curves: the equilibrium condition $\rho_{11}^{\mathrm{eq}}=\rho_{22}^{\mathrm{eq}}=1/2$ is applied with $\Gamma T_{1}=10$. Leakage is prevented and energy is trapped in the steady-state.
The leakage of internal energy is stopped in the steady-state by enforcing equilibrium conditions for the populations. To see how this happens we modify the time evolution equations for the two-atom populations obtained through the master equation by adding a relaxation term acting on a time-scale $T_{1}$ and a constant pump term, which we then define as $\rho_{jj}^{\mathrm{eq}}/T_{1}$, with $\rho_{jj}^{\mathrm{eq}}$ the equilibrium condition. We thus write
$\frac{d\rho_{11}}{dt}$ $=-\Gamma\left(1+F_{12}\right)\rho_{11}-\frac{\left(\rho_{11}-\rho_{11}^{\mathrm{eq}}\right)}{T_{1}}$ (33)
$\frac{d\rho_{22}}{dt}$ $=-\Gamma\left(1-F_{12}\right)\rho_{22}-\frac{\left(\rho_{22}-\rho_{22}^{\mathrm{eq}}\right)}{T_{1}}$ (34)
$\frac{d\rho_{33}}{dt}$ $=\Gamma\left(1+F_{12}\right)\rho_{11}+\Gamma\left(1-F_{12}\right)\rho_{22}-\frac{\left(\rho_{33}-\rho_{33}^{\mathrm{eq}}\right)}{T_{1}}.$ (35)
The steady-state solution of these equations yields
$\rho_{11}^{\mathrm{ss}}$ $=\frac{\rho_{11}^{\mathrm{eq}}}{\left(1+F_{12}\right)\Gamma T_{1}+1}\simeq\frac{\rho_{11}^{\mathrm{eq}}}{\left(1+F_{12}\right)\Gamma T_{1}}$ (36)
$\rho_{22}^{\mathrm{ss}}$ $=\frac{\rho_{22}^{\mathrm{eq}}}{\left(1-F_{12}\right)\Gamma T_{1}+1}\simeq\rho_{22}^{\mathrm{eq}}$ (37)
$\rho_{33}^{\mathrm{ss}}$ $=\left(1+F_{12}\right)\Gamma T_{1}\rho_{11}^{\mathrm{ss}}+\left(1-F_{12}\right)\Gamma T_{1}\rho_{22}^{\mathrm{ss}}+\rho_{33}^{\mathrm{eq}}$
$\simeq\rho_{11}^{\mathrm{eq}}+\left(1-F_{12}\right)\Gamma T_{1}\rho_{22}^{\mathrm{eq}}+\rho_{33}^{\mathrm{eq}},$ (38)
where the approximations apply whenever $\Gamma T_{1}\gg 1$ and $F_{12}\left(kz_{0}\right)\simeq 1$. Since internal energy trapping is due to non-zero steady-state values in the upper populations $\rho_{11}$ and $\rho_{22}$, we find that the presence of equilibrium conditions effectively causes this trapping to take place. One also readily verifies from equations (36)-(37) that no energy can be trapped when $T_{1}\rightarrow\infty$, which is the case shown with the black curves in Figure 2.
In Figure 2 we also show, using the colored curves and legend, the two-atom system response when $\rho_{11}^{\mathrm{eq}}=\rho_{22}^{\mathrm{eq}}=1/2$, $\rho_{33}^{\mathrm{eq}}=0$ and $\Gamma T_{1}=10$. The energy trapping in the two $m=0$ states is readily observed, with $\approx 0.5$ of the total energy remaining in $m=0$ states and a resulting energy trapping efficiency $\eta\simeq 0.46$.
Such analysis can be extended to cases involving a larger number of atoms. It will then generally be the case that the coherences $\rho_{ij}$, with $i\neq j$, will come into play (they do not affect our calculations for the two-atom populations). We therefore follow the usual practice of introducing a dephasing term scaling as $-\rho_{ij}/T_{2}$, with $T_{2}$ the corresponding time-scale, to account for external factors acting to reduce coherence in the system.
Figure 3: The master equation solution for a system consisting of two two-atom pairs located a distance $0.1\lambda$ apart; the two atoms forming one pair are still separated by $\Delta r=0.02\lambda$. Only the first two-atom pair is initially excited, equally in its fast and dark states. We set the equilibrium conditions for the populations of the four $m=-1$ states (i.e., |$\ket{aa}\ket{1,0}$, |$\ket{aa}\ket{0,0}$, etc.) to $1/4$ and that of the ground state to zero. The black curves show the time evolution of the relevant populations for $\Gamma T_{1}=\Gamma T_{2}=10$, while the red and blue curves show that of the ground state for $\Gamma T_{2}=1$ and $0.1$, respectively; $T_{1}$ is unchanged. Although the increased decoherence reduces the energy trapping efficiency from $\approx 0.6$ to $\approx 0.5$, energy trapping still occurs.
[…]
The effect of the dephasing time-scale is also depicted in the figure with the red and blue curves for the population of the |$\ket{aa}\ket{aa}$ state when $\Gamma T_{2}=1$ and $0.1$, respectively; $T_{1}$ is unchanged. It is found, as could be expected, that increased decoherence (i.e., a lower $T_{2}$) reduces the energy trapping efficiency, but it does not eradicate the effect. In fact, as can be inferred from the results, and a comparison with that from the single two-atom pair of Figure 2, decoherence tends to render the two pairs more uncorrelated.
Figure 4: Intensities (normalized to $\hbar\omega_{0}\Gamma$) for the single-atom, and two- and four-atom cases of Figures 2 and 3 when $\Gamma T_{1}=\Gamma T_{2}=10$. While equilibrium conditions have enable the trapping of energy in the long run, it also brings a non-zero steady-state intensity in all cases. Subradiance and energy trapping are apparent from the narrower profiles of the transient intensities for the two- and four-atom systems over the single-atom case.
Figure 4 shows the intensities (normalized to $\hbar\omega_{0}\Gamma$) for the single-atom, and the two- and four-atom cases of Figures 2 and 3 when $\Gamma T_{1}=\Gamma T_{2}=10$. While equilibrium conditions have enabled the trapping of energy in the long run, it also brings a non-zero steady-state intensity in all cases. This is not surprising since, as discussed above, equilibrium forces the establishing of non-zero steady-state populations. The systems then radiate in this regime at intensities proportional to the sum of the products of the populations and their transitions rates (see equation 19). In the figure all three systems radiate at approximately similar steady-states intensities because of the relatively high value of $\Gamma T_{1}=10$; lower values would generally differentiate between them.
Subradiance and energy trapping are also apparent in Figure 4 from the shorter time-scales (i.e., narrower profiles) of the transient intensities for the two- and four-atom systems over the single-atom case. But it should be kept in mind that these results all rest on the fact that $k\Delta r\ll 1$ for the close pairs of atoms. These are not common conditions in the interstellar medium. It would thus appear that a many-atom system would behave in the same manner as a single-atom whenever $k\Delta r\gg 1$.
There is, however, another way toward efficient energy trapping. As was observed in Sec. II.1, environmental conditions can sometimes favor the realization of specific radiation modes. In such cases, radiation can become coherent and transition rates can approach those of the small sample systems (see equations 11-14 or 24-27). We now consider this possibility.
IV Superabsorption, coherence and subradiance
Extending the analysis of Sec. III would quickly become prohibitive for even a moderate number of atoms. Furthermore, we have limited ourselves to the emission of a single photon from the $m=-n/2+1$ states to the $m=-n/2$ ground state. Realistic conditions would require initiating the system at higher $m$ values. In order to generalize our analysis we now move to the framework of the Maxwell-Bloch equations (MBE).
In this section, we study the propagation of an incident radiation field on a slab composed of an atomic gas, similar to what is often done in elementary radiative transfer analysis in astrophysics [25]. This will allow us to compare the outcome of our analysis with well-known results. The MBE are perfectly suited for this problem, being one-dimensional in nature. We will also investigate effects the incident field will have on the radiation emanating from the gas itself.
Assuming propagation along the $z$-axis, the MBE [20, 6, 13] for a magnetic dipole transition are [7]
$\frac{\partial n^{\prime}}{\partial\tau}$ $=\frac{i}{\hbar}\left(M^{+}B^{+}-B^{-}M^{-}\right)-\frac{n^{\prime}-n^{\prime}_{\mathrm{eq}}}{T_{1}}$ (39)
$\frac{\partial M^{+}}{\partial\tau}$ $=\frac{2i\mu^{2}}{\hbar}B^{-}n^{\prime}-\frac{M^{+}}{T_{2}}$ (40)
$\frac{\partial B^{+}}{\partial z}$ $=\frac{i\mu_{0}\omega_{0}}{2c}M^{-},$ (41)
where $n^{\prime}=\left(n_{b}-n_{a}\right)/2$ with $n_{a}$ and $n_{b}$ the densities of atoms in the lower and upper states, respectively, and $M^{+}$ and $B^{+}$ the amplitudes of the molecular magnetization and the magnetic induction field. The temporal derivatives are in relation to the retarded time $\tau=t-z/c$. The ‘slowly varying envelope approximation’ is used with the magnetization and magnetic field vectors given by
$\mathbf{M}^{\pm}\left(z,\tau\right)$ $=M^{\pm}\left(z,\tau\right)e^{\pm i\omega_{0}\tau}\bm{\epsilon}_{\mu}$ (42)
$\mathbf{B^{\pm}}\left(z,\tau\right)$ $=B^{\pm}\left(z,\tau\right)e^{\mp i\omega_{0}\tau}\bm{\epsilon}_{\mu}.$ (43)
As in Sec. III, we added relaxation and dephasing terms with corresponding time-scales $T_{1}$ and $T_{2}$. We can relate this framework to that used in previous sections through the correspondences
$n^{\prime}$ $\rightarrow\frac{1}{\delta\mathcal{V}}\sum_{j\in\delta\mathcal{V}}\hat{R}_{j}^{3}$ (44)
$M^{+}$ $\rightarrow\frac{\mu}{\delta\mathcal{V}}\sum_{j\in\delta\mathcal{V}}\hat{R}_{j}^{+}e^{ikz_{j}},$ (45)
at a given position $z$ with $\delta\mathcal{V}$ the volume occupied by the gas in the vicinity. In what follows, we will consider $n_{-}\equiv 2n^{\prime}=n_{b}-n_{a}$ instead of $n^{\prime}$. Furthermore, we will mainly be concerned with conditions of thermal equilibrium or, more generally, situations where $n_{b}<n_{a}$. We thus have, unless otherwise indicated, $n_{-}<0$.
With the incident radiation field propagating along the $z$-axis, as expressed in equation (43) and in the definition of the retarded time, we have an example of situations where environmental or external conditions favor one radiation mode over others. Importantly, the interaction between the incident field and the gas will initially elicit superabsorption and favor coherence in subsequent transitions in the gas through angular correlation among photons, as discussed at the end of Sec. II.1. It is therefore expected that the radiation field emanating from, and propagating within, the gas itself will also take the form of equation (43).
Hence, we will subject the system to two independent stimuli: i) the aforementioned incident magnetic field $B_{0}\equiv B^{+}\left(0,\tau\right)$, which effectively serves as a boundary condition at the input $z=0$ of the system, and ii) internal quantum fluctuations throughout the system at $\tau=0$ modeled as an initial magnetization $M^{+}\left(z,0\right)=n_{-}^{\mathrm{eq}}\mu\,\theta_{0}/2$, where $\theta_{0}=\sqrt{2N}/\left|N_{-}^{\mathrm{eq}}\right|$ is the initial “rising” angle [20] with $N=N_{a}+N_{b}$ the total number of atom ($N_{a}$ and $N_{b}$ in the lower and upper states, respectively) and $N_{-}^{\mathrm{eq}}=N_{b}^{\mathrm{eq}}-N_{a}^{\mathrm{eq}}$ (see below and Appendix B). Unlike for our discussion of Dicke’s work in Sec. II, we express numbers of atoms using capitalized letters from now on to avoid potential confusion with number densities.
Finally, since phase coherence is expected for the radiation field and that this cannot take place over an arbitrarily large volume, we will enforce a Fresnel number of unity and focus our analysis on a cylinder of cross-section area $A=\lambda L$, with $L$ the length of the cylinder and $\lambda$ the wavelength of radiation [20]. We then have for the relevant numbers of atoms $N_{a}=n_{a}AL$, etc. Our slab is therefore composed of a very large number of independent cylinders.
IV.1 Response to incident radiation
IV.1.1 Linear regime
The general solution of the MBE requires numerical computations. It is, however, possible to provide analytical solutions in the linear regime (i.e., weak radiation intensity) where $n_{-}=n_{-}^{\mathrm{eq}}$. We show in Appendix A that the magnetic field resulting from the response of the gas to an incident field of amplitude $B_{0}$ is
[…]
$\left.+\frac{\alpha}{2}e^{-\tau/T_{2}}\left[e^{-\alpha z/2}\star J_{0}\left(2\sqrt{\frac{z\tau}{LT_{R}}}\right)\right]\right\},$ (46)
where $J_{p}\left(x\right)$ is the Bessel function of the first kind and order $p$, while $\star$’ stands for a (spatial) convolution. The two terms on the right-hand side consist of the steady-state and transient responses, respectively, with the absorption coefficient generally defined as (i.e., for any density $n_{-}$)
$\alpha$ $=\frac{\mu_{0}\omega_{0}\mu^{2}T_{2}\left|n_{-}\right|}{\hbar c}$
$=\frac{2T_{2}}{LT_{R}}.$ (47)
In the last equation we have introduced the superabsorption time-scale (defined at equilibrium)
$T_{R}=\frac{8\pi}{3\lambda^{2}\left|n_{-}^{\mathrm{eq}}\right|L\Gamma}$ (48)
which has exactly the same form as the superradiance time-scale, except for the fact that in our case $n_{-}^{\mathrm{eq}}<0$ [21, 20, 6].
We will soon be interested in estimating the duration of the transient response of the system in the limit where $T_{2}\ll T_{R}$. As a first step, expanding the Bessel function for $z\tau\ll LT_{R}$ in equation (46) allows us to approximate
$B^{+}\left(z,\tau\right)\simeq B_{0}$ $\left[e^{-\alpha z/2}\left(1-e^{-\tau/T_{2}}e^{-2\sqrt{z\tau/LT_{R}}}\right)\right.$
$\left.+e^{-\tau/T_{2}}e^{-2\sqrt{z\tau/LT_{R}}}\right].$ (49)
If we further restrict ourselves to $T_{2}\ll T_{R}$, then expanding the square root about $zT_{2}/LT_{R}$ in equation (49) yields
$e^{-\tau/T_{2}}e^{-2\sqrt{z\tau/LT_{R}}}\approx e^{-\tau\left(1+\sqrt{zT_{2}/LT_{R}}\right)/T_{2}}$ (50)
and an evolution time-scale
$T_{\mathrm{tr}}\approx T_{2}\left(1+\sqrt{\frac{zT_{2}}{LT_{R}}}\right)^{-1}.$ (51)
The transient signal thus dies out on a time-scale of the order of $T_{\mathrm{tr}}$. This relation will be useful when studying the steady-state response in intensity in more details. This is what we now tackle before applying the MBE to the 21 cm transition of atomic hydrogen in an astrophysical context.
IV.1.2 Steady-state solution
It is also instructive to study the steady-state response for more general conditions, i.e., without limiting ourselves to the linear regime. Setting the time derivatives in equation (39)-(40) to zero, we readily find
$n_{-}$ $=\frac{n_{-}^{\mathrm{eq}}}{1+I/I_{\mathrm{sat}}}$ (52)
$\frac{dI}{dz}$ $=-\alpha I,$ (53)
where the intensity $I=cB^{+}B^{-}/2\mu_{0}$, and the saturation intensity
$I_{\mathrm{sat}}=\frac{\hbar^{2}c}{8\mu_{0}\mu^{2}T_{1}T_{2}}.$ (54)
The steady-state solution leads to relations (i.e., equations 47 and 52-54) that have exactly the same form as the maser equations with, again, the difference that $n_{-}<0$ [3, 6]. Accordingly, two regimes are identified for the intensity.
In the linear regime, when $I\ll I_{\mathrm{sat}}$, we find that $n_{-}\left(z,\tau\right)=n_{-}^{\mathrm{eq}}$ and
$I\left(z\right)=I_{0}e^{-\alpha z}$ (55)
with $I_{0}=c\left|B_{0}\right|^{2}/2\mu_{0}$ and $\alpha$ a constant (i.e., with $n_{-}=n_{-}^{\mathrm{eq}}$ in equation 47). This solution is the same as found in the steady-state component of equation (46), and is of a similar form as Beer’s law when $\alpha$ replaced by [26]
$\alpha_{\mathrm{B}}=\frac{3}{4}\left|n_{-}^{\mathrm{eq}}\right|\lambda^{2}\Gamma\phi\left(\omega\right)$ (56)
[…]
A comparison of equations (47) and (56) reveals that
$\alpha=\alpha_{\mathrm{B}}T_{2}\delta\omega$ (57)
with $\delta\omega$ the spectral extent of $B^{+}\left(z,\tau\right)$ as it propagates through the gas. This quantity is the reciprocal of the transient response time-scale given in equation (51). We then find
$T_{2}\delta\omega\approx 1+\sqrt{\frac{zT_{2}}{LT_{R}}}$ (58)
and
$\alpha\approx\alpha_{\mathrm{B}}\left(1+\sqrt{\frac{zT_{2}}{LT_{R}}}\right).$ (59)
for $T_{2}\ll T_{R}$.
We uncover the important result that Beer’s law is only truly valid when $T_{2}\rightarrow 0$, i.e., in the limit of complete decoherence. In the opposite limit when $T_{2}\gg T_{R}$, we have $\delta\omega\sim T_{R}^{-1}$, $\alpha=\alpha_{\mathrm{B}}T_{2}/T_{R}$ and find, just as importantly, that the density of the gas determined with Beer’s law can be overestimated when coherence is present.
The saturated regime when $I\gg I_{\mathrm{sat}}$ will be considered in more details in a future publication. But we note its solution
$I\left(z\right)=I_{0}-\frac{\left|n_{-}^{\mathrm{eq}}\right|\hbar\omega_{0}}{8T_{1}}z$ (60)
for $z<z_{\mathrm{sat}}$, where $z_{\mathrm{sat}}$ is the location where $I=I_{\mathrm{sat}}$. This regime is entirely neglected when Beer’s law is solely considered.
Figure 5: Steady-state intensity as a function of position along the optical path in the gas. This result was obtained with the MBE (i.e., equations 39-41). The linear and saturated regimes are indicated, as well as the saturation intensity $I_{\mathrm{sat}}$.
The steady-state solution obtained with the MBE (i.e., equations 39-41) is shown in Figure 5 with the intensity plotted as a function of the position within the gas. The linear and saturated regimes are indicated, as well as the saturation intensity $I_{\mathrm{sat}}$. The shape of the intensity curve is reminiscing, and practically a mirror image, of the one obtained for masers when $n_{-}>0$ (see Fig. 4 in [6]).
Figure 6: Left: Temporal evolutions of the radiation intensity (top) and magnetic induction field (bottom) of the 21 cm line for an atomic hydrogen gas at a temperature of $10$ K and system length $L\simeq 1.5$ pc. We set $T_{1}=T_{2}\simeq 16\,T_{R}$ and the intensity is normalized to the saturation intensity $I_{\mathrm{sat}}$ given in equation (54). In the bottom panel, the incident magnetic field’s amplitude is shown in cyan (using the vertical scale on the right), while a broken curve in red corresponding to equation (46) is plotted on top of that obtained from numerical computations (in black) using the MBE (i.e., equations 39-41). Right: Temporal evolution of the density $n_{-}$ in black with the pump $\Lambda_{n}=n_{-}^{\mathrm{eq}}/T_{1}$ in cyan (using the vertical scale on the right; top) and the magnetization (bottom). The density is limited to the spectral extent of the signal (i.e., $T_{R}^{-1}$) and corresponds to a total density of $1\,\mathrm{cm^{-3}}$ for a gas at a temperature of $10$ K. The mean distance between neighboring cooperating atoms is $\Delta r\approx 670\lambda$ and $T_{R}=1.5\times 10^{8}\,\mathrm{s}\simeq 4.3\times 10^{-7}/\Gamma$.
IV.1.3 Linear regime response – The 21 cm line
We now consider a slab of atomic hydrogen gas at thermal equilibrium with a density of $1\,\mathrm{cm^{-3}}$ and a temperature $T_{\mathrm{kin}}=10$ K. This approximately corresponds to typical mass densities measured in Dark Matter halos [27, 28], while we assume the gas to be cold. Thermal motions bring an average time-scale between collisions of about $2.5\times 10^{9}$ s, which we ascribe to $T_{2}$. Since we are considering radiation from the 21 cm line, which has an excitation temperature $T_{\mathrm{21\,cm}}\simeq 0.068\,\mathrm{K}\ll 10$ K, we set, for simplicity, $T_{1}\approx T_{2}=2.5\times 10^{9}$ s (i.e., we assume that inelastic and elastic collisions happen at similar rates and are the sole sources of relaxation and dephasing in the gas, respectively).
Because we expect the linear regime response of the system to happen on the superabsorption time-scale $T_{R}$ (see equation 46) it follows that only a fraction of the atomic population can act cooperatively in the process. That is, given the spectral breadth $\Delta\omega\sim\omega_{0}\Delta v/c$ associated with thermal motions (of characteristic speed $\sim\Delta v$) the density entering our equations is $1\,\mathrm{cm^{-3}}/T_{R}\Delta\omega$. In accordance with the so-called Arecchi-Courtens condition [29], we fix the thickness of the slab to $L=cT_{R}$, which corresponds to the maximum length of a coherent region when the excitation happens simultaneously over the whole sample (see Sec. V below). Using equation (48) and the Boltzmann equation for the population of energy levels we arrive at $n_{-}^{\mathrm{eq}}=9.6\times 10^{-15}\,\mathrm{cm^{-3}}$, $L=4.5\times 10^{18}\,\mathrm{cm}\simeq 1.5$ pc and $T_{R}=1.5\times 10^{8}\,\mathrm{s}=4.3\times 10^{-7}/\Gamma^{-1}$, with $\Gamma\simeq 2.9\times 10^{-15}\,\mathrm{s^{-1}}$ for the 21 cm line. We therefore set a constant “pump” $\Lambda_{n}=n_{-}^{\mathrm{eq}}/T_{1}$ to enforce thermal equilibrium and keep $n_{-}=n_{-}^{\mathrm{eq}}$ at all times (see equation 39). We also note that the mean distance between neighboring cooperating atoms is $\Delta r\approx 670\lambda$ and $T_{2}\simeq 16T_{R}$, which implies that coherence can be sustained in the system under these conditions.
The response of the system to a constant incident radiation field $B^{+}\left(0,\tau\right)\equiv B_{0}=1\times 10^{-23}$ T, as numerically computed with the MBE (i.e., equations 39-41), is shown in Figure 6. This corresponds to an incident intensity approximately five orders of magnitude lower than the saturation intensity $I_{\mathrm{sat}}$, which ensures that we are well into the linear regime. The precise intensity level is irrelevant to our discussion, as long as we remain in the linear regime.
The left panel of Figure 6 shows the temporal evolutions of the radiation intensity (normalized to $I_{\mathrm{sat}}$; top) and the magnetic field (bottom). A broken curve in red corresponding to equation (46) is plotted on top of that obtained from numerical computations using the MBE for the magnetic field (in black); there is perfect agreement between the two. The (constant) density $n_{-}$ (in black) and pump $\Lambda_{n}$ (in cyan; using the vertical scale on the right; top) and the magnetization (in units of Bohr magneton $\mu_{\mathrm{B}}$ per cubic centimeter; bottom) are displayed in the right panel.
It is interesting to note that at $\tau=0$ the incident magnetic field is transmitted unattenuated at $z=L$ (or any other position) no matter how large $L$ is. Thereafter, superabsorption defines the transient response to greatly reduce the intensity on a time-scale $\approx T_{R}$ and in a oscillatory manner reminiscing of superradiance systems. According to our previous discussions, the steady-state regime settles afterwards to an intensity diminished by a factor $e^{-\alpha L}$, where $\alpha L=2T_{2}/T_{R}\simeq 32$ from equation (47). Such signal would therefore be detected as a deep and narrow absorption line ($\approx 0.5\,\mathrm{km\,s^{-1}}$ in width) against the background radiation.
Finally, we note that the behavior of such a system is mainly determined by the ratio of the dephasing time-scale $T_{2}$ to that for subabsorption $T_{R}$, which in turn is completely dependent on the column density $\left|n_{-}^{\mathrm{eq}}\right|L$. It follows that in instances where $T_{1}$ and $T_{2}$ are not set by collisions, changes in $n_{-}^{\mathrm{eq}}$ can be offset by opposite ones in $L$ to preserve the system’s response (and vice-versa).
IV.2 Response to initial fluctuations
As shown in Appendices A and B, the linear regime response of the system to initial internal fluctuations $M^{+}\left(z,0\right)=n_{-}^{\mathrm{eq}}\mu\theta_{0}/2$, with $\theta_{0}=\sqrt{2N}/\left|N_{-}^{\mathrm{eq}}\right|$, is expressed as
$B^{+}\left(z,\tau\right)=-i\frac{\hbar\theta_{0}}{2\mu T_{R}}e^{-\tau/T_{2}}\frac{z}{L}\sqrt{\frac{LT_{R}}{z\tau}}J_{1}\left(2\sqrt{\frac{z\tau}{LT_{R}}}\right).$ (61)
Studying the behavior of $J_{1}\left(2\sqrt{x}\right)/\sqrt{x}$ reveals that the amplitude of the field corresponding to the initial fluctuations in magnetization is transmitted without amplification/absorption to $z=L$ at $\tau=0$. This response is linear in $z$ and therefore coherent (i.e., the intensity is proportional to $z^{2}$). At a given position, the field (and the corresponding intensity) displays a weak oscillatory temporal behavior with a period scaling with $T_{R}$, but damped on the dephasing time-scale $T_{2}$.
We show in Figure 7 the response to these internal fluctuations for the same system as in Sec. IV.1.3 when studying the propagation of an incident radiation field. These results were obtained by integrating the MBE (i.e., equations 39-41). The left panel of the figure presents the temporal evolution of the radiation intensity (top) and magnetic field (bottom). The intensity is normalized to that expected from a gas radiating non-coherently with $N_{b}$ excited atoms in these conditions [7, 8]
$I_{\mathrm{nc}}$ $=N_{b}\frac{\hbar\omega_{0}\Gamma}{A}\frac{\phi_{\mathrm{D}}}{8\pi/3}$
$=\frac{n_{b}}{\left|n_{-}^{\mathrm{eq}}\right|}\frac{\hbar\omega_{0}}{AT_{R}},$ (62)
where $\phi_{\mathrm{D}}=\lambda^{2}/A$ is the diffraction solid-angle at the end of our (cylindrical) system. The right panel displays the temporal evolution of the density $n_{-}$ in black with the pump $\Lambda_{n}=n_{-}^{\mathrm{eq}}/T_{1}$ in cyan (using the vertical scale on the right; top) and the magnetization (bottom).
Figure 7: Left: Temporal evolution of the radiation intensity (top) and magnetic field (bottom) of the 21 cm line for the same atomic hydrogen gas as in Figure 6. That is, the gas temperature is $10$ K, the length $L=1.5$ pc, $T_{1}=T_{2}\simeq 16\,T_{R}$, with $T_{R}=1.5\times 10^{8}\,\mathrm{s}\simeq 4.3\times 10^{-7}/\Gamma$, and $\Delta r\approx 670\lambda$. The intensity is normalized to that expected from a gas radiating non-coherently under the same conditions. A broken curve in red corresponding to equation (61) is plotted on top of that obtained from numerical computations (in black) using the MBE (i.e., equations 39-41). Right: Temporal evolution of the density $n_{-}$ in black with the pump $\Lambda_{n}=n_{-}^{\mathrm{eq}}/T_{1}$ in cyan (using the vertical scale on the right; top) and the magnetization (bottom).
The curve for the normalized intensity is reminiscent of those found in Figure 4 for the simple two- and four-atom systems discussed in Sec. III using the master equation. Here, however, we have $N\sim 10^{27}$ cooperating atoms in total contained in the system ($\left|N_{-}^{\mathrm{eq}}\right|\sim 10^{24}$), which in comparison leads to an extreme level of subradiance. More precisely, at $\tau=0$ the system’s intensity is at a level similar to what is expected from a non-coherent gas (i.e., $I=0.25\,I_{\mathrm{nc}}$), but it only radiates for a duration $\approx T_{R}$. To that effect, we should emphasize that $\Gamma T_{R}\simeq 4.3\times 10^{-7}$. In comparison, for the 21 cm line the non-coherent transient intensity for $N_{b}$ non-cooperating atoms would last on the order of $\Gamma^{-1}\sim 10^{14}\,\mathrm{s}\sim 10^{7}$ yr, whereas $T_{R}\sim 5$ yr. This implies that almost the totality of the available energy is locked within the atomic population, which is almost entirely in dark states since the fast states are quickly depleted during the transient regime.
Unlike the two- and four-atom systems in Figure 4, the steady-state intensity is zero despite having a significant fraction (almost half) of the atoms in the excited state. The combination of the widespread entanglement of atoms and coherence in our system suppresses the intensity in the steady-state regime. We therefore find that the gas becomes completely dark after a very short time.
Finally, we note the fact that, from equation (61), decreasing the dephasing time-scale $T_{2}$ in the system does not stop it from becoming dark. If anything, it reduces the time needed for this phenomenon to take place. This is because, when using the MBE, coherence is locked in by the fact that one radiation mode is favored over all others (see equation 43). The main effect in decreasing $T_{2}$ is therefore limited to increasing the dampening of the field in equation (61).
V Discussion – Radiation anisotropy, global atomic entanglement and Dark Matter
In the previous sections we have discussed the requirements leading to the suppression of radiation emanating from a gas in the steady-state regime. We found that the entanglement of a large number of atoms and the establishing of coherence are the main conditions to be fulfilled for this effect to take place.
If entanglement naturally results from the interaction between the atoms composing the gas and the ambient electromagnetic field, coherence can only happen under certain conditions. As reviewed in Secs. II and III, close proximity between atoms (i.e., separations much shorter than $\lambda$; the small sample limit) ensures coherence for all radiation modes. The manifestation of this coherence will take different forms depending on the initial conditions. For example, superradiance will ensue when the gas is initially hosting a population inversion. Otherwise, at thermal equilibrium, for example, the initial occupation of slow and dark states leads to subradiance. Importantly, dark states are responsible for the trapping of energy and the suppression of radiation in the steady-state. However, close proximity between cooperating radiators is unlikely to be common in astrophysical media and the trapping of radiation is negated under such conditions, as energy escapes through non-coherent radiation modes.
Equilibrium conditions set the population of slow and dark states, and will naturally stop their gradual decay in the steady-state. These features render equilibrium essential to achieving significant levels of subradiance. However, this comes at the cost of a non-zero steady-state intensity (see Figure 4). Furthermore, equilibrium conditions do not improve the energy trapping efficiency when the atoms are not in close proximity, since it cannot ensure that coherence in the radiation field is established.
As mentioned at the end of Sec. II.1, close proximity between atoms is not the only way to achieve energy trapping. As is evident from the transition rates for the two-atom system (see equations 13-14 and 26-27), no matter how large the separation between atoms, coherence can be achieved through angular correlation between emitted/absorbed photons. This will be realized whenever one mode of radiation (or a few) is favored over the others. This can happen whenever some form of anisotropy is present, whether in the geometry of the system (e.g., elongated structures) or the ambient physical conditions (e.g., accentuated velocity coherence along a well-defined direction). Although these forms of anisotropy are also applicable to the problem of the atomic hydrogen 21 cm line analyzed in Sec. IV, we found that another scenario is likely to be efficient in that respect.
In our analysis of the 21 cm line in Sec. IV, we purposely chose a gas density that is in line with measurements in galactic Dark Matter halos. For the matter contained in these environments, the main source of anisotropy comes in the form of the radiation field emanating from the host galaxy. As then mentioned, the incidence of this field will effectively select one mode for the propagation of radiation within the halo. It is expected that superabsorption of the incident field will elicit coherence in subsequent transitions in the gas through angular correlation among photons (see Sec. II.1). As a consequence, extreme subradiance, leading to a complete darkening of the gas, takes place within a few years even at a mean separation $\Delta r\approx 670\lambda$ between closest neighbors. As far as emission properties are concerned, this atomic hydrogen gas is invisible and effectively behaves like Dark Matter. This darkening effect may therefore possibly account for some of the Dark Matter known to exist in these environments.
However, is the suppression of radiation in a given transition through subradiance sufficient for the corresponding atomic species to be considered dark? In other words, what about all the other transitions pertaining to this atom?
It is important to note that entanglement is not a property that is unique to a specific transition. That is, it is always possible to use the symmetry of the Hamiltonian to determine proper states for the problem at hand, such as those shown in equations (7)-(9) for the two-atom system. The forms of these symmetry-adapted states are the same for all energy levels. For example, if equations (7)-(9) apply to the states pertaining to the 21 cm transition for two entangled atoms, then similar looking states will also apply for, say, the Lyman-$\alpha$ or other hydrogen lines. In other words, entanglement is a global property of the atom. Coherence, however, is not and needs to be assessed on a transition by transition basis.
For our analysis of the 21 cm line for a gas at a density $n_{-}^{\mathrm{eq}}=1\,\mathrm{cm^{-3}}$ and temperature $T_{\mathrm{kin}}=10$ K these considerations have no bearing on the outcome since all other lines (e.g., Lyman-$\alpha$) will remain essentially unexcited. The gas as a whole will thus appear dark in the steady-state on the account of the 21 cm line alone. While we only studied this specific example, similar conclusions may be reached when considering other environments with differing gas densities and temperatures. Moreover, we should also note that this darkening effect is potentially not limited to the hydrogen atom. For example, the hydroxyl molecule (OH) has several lines at radio wavelengths that could be considered. It also follows that this darkening effect, which also relies on the presence of anisotropic radiation, should be closely correlated with the presence of “luminous” matter, as is the case for Dark Matter [28].
We also studied the propagation of an incident radiation field at 21 cm in the same atomic hydrogen gas. Interestingly, we found that Beer’s law, commonly used to quantify the attenuation in intensity as a function of position, is an approximation that only applies under complete decoherence (i.e., in the limit $T_{2}\rightarrow 0$) in the linear regime. It follows that the amount of attenuation could exceed expected levels in situations where coherence exists in the gas. Furthermore, the existence of the saturated regime is completely missed when only Beer’s law is considered.
We are then led to a picture where our “dark” hydrogen gas (in emission) would also appear more opaque (or more optically thick) in its absorption lines against the background source responsible for the incident radiation. However, it would otherwise be completely transparent to radiation at wavelengths alien to the hydrogen atom, as expected from Dark Matter. For a cold gas as the one considered in our analysis, these lines would appear as narrow absorption features. Such spectral features are commonly detected in observations of the interstellar medium [30], the Galactic halo [31], the Galactic disk [32] and external galaxies [33].
Another interesting and important property of our darkened atomic hydrogen gas is discussed in Appendix C. There, we calculate the cross-section for collisions between two two-atom systems, both being entangled as discussed in Sec. II.1. Interestingly, it is found that this collision cross-section vanishes because of destructive interference due to the entangled nature of the states the systems find themselves in after the transient regime. Although this analysis needs to be generalized and formalized, this result is consistent with astronomical observations of the so-called bullet cluster and similar objects [34]. These observations put a constraint on the ratio of the collision cross-section to the mass of a purported Dark Matter (point) particle to $\sigma/M\lesssim 1\,\mathrm{cm^{2}\,g^{-1}}$ [28]. Although the corresponding figure for atomic hydrogen is $\sim 10^{9}\,\mathrm{cm^{2}\,g^{-1}}$, it does not apply in our case. Our analysis does not suggest that atomic hydrogen is a Dark Matter particle, but rather that the cooperative behavior of a large number of entangled hydrogen atoms is consistent with what is expected for Dark Matter under the conditions we considered. While individual atoms may still collide, entangled systems of atoms will be essentially collision-less.
Finally, we should make a few observations concerning our model based on the MBE, as presented in Sec. IV. First, we used the Arecchi-Courtens condition to set the length scale of a coherent entity within the gas [29]. In superradiance systems, for example, this is applicable to cases where the pump responsible for the population inversion is transverse, i.e., it affects the gas simultaneously at all positions. When this happens, regions separated by a distance $L>cT_{R}$ will evolve on a time-scale ($T_{R}$) shorter than that needed for their interaction ($L/c$) and, therefore, cannot work cooperatively. In our case, thermal equilibrium is set by the pump and one would therefore expect the Arecchi-Courtens condition to apply. However, the favoring of a single radiation mode, through which the photon angular correlation and coherence are set, can only happen when upstream photons reach atoms located further down along the optical path. This is akin to the effect of a longitudinal pump, in which case the Arecchi-Courtens condition would not apply [20]. This would then relax any restriction on the coherent length in the gas and allow for lower values of $T_{R}$, and potentially broaden the applicability and the strength of the darkening effect (see below).
On the other hand, the use of the MBE renders it difficult to investigate situations where coherence and the darkening effect cannot be established. This is because the selection of a single radiation mode is inherent in the model (see equations 42-43). The only manner coherence can be impeded is, therefore, through the lowering of the dephasing time-scale $T_{2}$. But as we stated at the end of Sec. IV.2, for subradiance the impact of a lower $T_{2}$ is limited to the dampening of the signal. Along the same line of thoughts, while we set $T_{2}$ to the collision time-scale for simplicity in our analysis, other factors (e.g., interaction with non-coherent modes from external radiation) could reduce its value and further increase dephasing in the system.
This limited effect of dephasing in our system stems from the fact our model is one-dimensional, while realistic systems are three-dimensional in nature. For example, the darkening effect clearly cannot proceed for a more-or-less spherical cloud of gas when the ambient field is not sufficiently anisotropic to ensure the selection of one (or a few) radiation mode. To numerically test this scenario, one would need a three-dimensional generalization of the MBE to allow for differing boundary conditions (i.e., different types of incident field) and geometries for the cloud of gas. This would also permit the investigation of the effect of dephasing and decoherence when a large number of radiation modes are available. This may reveal the lack of a darkening, or some limitation of it, for sufficiently low values of $T_{2}$.
These considerations further emphasize the relevance of our earlier discussion on the applicability of the Arecchi-Courtens condition. This may become consequential for the assessment of the importance of the darkening effect, mainly because of the impact of the kinetic temperature on the time- and length-scales involved. For example, when $T_{\mathrm{kin}}=50$ K we find that the Arecchi-Courtens condition leads to $L\approx 16$ pc, $T_{R}\approx 50$ yr and $T_{2}\approx 0.6\,T_{R}$, while $L\approx 46$ pc, $T_{R}\approx 150$ yr and $T_{2}\approx 0.16\,T_{R}$ for $T_{\mathrm{kin}}=100$ K. The values obtained for $L$ are reasonable considering the size of Dark Matter halos and those for $T_{R}$ are negligible on astrophysical time-scales, but the decrease in $T_{2}$ relative to $T_{R}$ implies a corresponding increase in decoherence. If the Arecchi-Courtens condition were not to apply for the reasons previously stated, then the level of coherence could be re-established by simply increasing $L$ (with a corresponding decrease in $T_{R}$; see equation 48). If, on the other hand, it does apply, then it becomes important to verify our MBE-based predictions with three-dimensional models.
Data availability
The data that support the findings of this article are available from the authors upon reasonable request.
Acknowledgements.
We thank J. Cami for his helpful comments. M.H.’s research is funded through the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant RGPIN-2024-05242 and the Western Strategic Support for Research Accelerator Success. F.R.’s research is supported by the NSERC Discovery Grant RGPIN-2024-06346. M.H. and F.R. are grateful for the hospitality of Perimeter Institute where part of this work was carried out.
Authors contributions
M.H. conceptualized the work. Both authors contributed equally to the analysis and the writing of the paper.
Appendix A Linear solutions of the Maxwell-Bloch equations
In the linear regime when $I\ll I_{\mathrm{sat}}$, the magnetic field and the polarization are small and we expect $\partial n_{-}/\partial\tau\approx 0$, implying that $n_{-}=n_{-}^{\mathrm{eq}}$ is constant. Equations (40)-(41) then reduce to a simple system of linear differential equations. These equations can be solved using Laplace transforms, with the $z\leftrightarrow u$ and $\tau\leftrightarrow s$ correspondences, leading to
$B_{1}^{+}\left(u,s\right)=B_{0}\frac{s+1/T_{2}}{us\left(s+1/T_{2}+1/uLT_{R}\right)}$ (63)
when the system is subjected to an incident magnetic field of amplitude $B_{0}\equiv B^{+}\left(z=0,\tau\right)$ and
$B_{2}^{+}\left(u,s\right)=i\frac{n_{-}^{\mathrm{eq}}\mu_{0}\omega_{0}\mu\theta_{0}}{4c}\frac{1}{u^{2}\left(s+1/T_{2}+1/uLT_{R}\right)}$ (64)
when responding to initial internal magnetization fluctuations $M^{+}\left(z,0\right)=n_{-}^{\mathrm{eq}}\mu\theta_{0}/2$. These two solutions are uncorrelated since $\theta_{0}$ is a random process.
Effecting a first inverse Laplace transform to recover the retarded time $\tau$ and then a second to recover $z$, using
$\mathcal{L}^{-1}\left\{\frac{e^{-a/u}}{u^{n}}\right\}=\left(\frac{z}{a}\right)^{\frac{n-1}{2}}J_{n-1}\left(2\sqrt{az}\right),$ (65)
with $J_{p}\left(x\right)$ the Bessel function of the first kind and order $p$ [35], we find
[…]
$\left.+\frac{\alpha}{2}e^{-\tau/T_{2}}\left[e^{-\alpha z/2}\star J_{0}\left(2\sqrt{\frac{z\tau}{LT_{R}}}\right)\right]\right\}$ (66)
$B_{2}^{+}\left(z,\tau\right)$ $=-i\frac{\hbar\theta_{0}}{2\mu T_{R}}e^{-\tau/T_{2}}\frac{z}{L}\sqrt{\frac{LT_{R}}{z\tau}}J_{1}\left(2\sqrt{\frac{z\tau}{LT_{R}}}\right)$ (67)
with $\alpha$ and $T_{R}$ defined in equations (47)-(48). In equation (66) $\star$’ stands for a spatial convolution.
Appendix B Initial Bloch angle and magnetization
We require initial conditions for the integration of the MBE in equations (39)-(41). While $n_{-}^{\prime}\left(z,0\right)=n_{-}^{\mathrm{eq}}/2$, it is clear from an inspection of the MBE that in the absence of an incident signal (i.e., when $B^{+}\left(0,\tau\right)=0$) we have $B^{+}\left(z,\tau\right)=M^{+}\left(z,\tau\right)=0$ if $M^{+}\left(z,0\right)=0$. In reality, internal quantum fluctuations modeled through the initial Bloch angle $\theta\left(0\right)$ (see below) will cause a corresponding magnetization that will start the response from the gas. Here, we closely follow the treatment developed for superradiance by Gross and Haroche [20] but adapt it to our problem.
With $\hat{R}^{\pm}=\sum_{j}\hat{R}^{\pm}_{j}$, we start by noting that [36]
$\hat{R}^{\pm}\Ket{r,m}=\sqrt{\left(r\mp m\right)\left(r\pm m+1\right)}\Ket{r,m\pm 1},$ (68)
which in turn leads to
$\left(\hat{R}^{-}\right)^{p}$ |$\Ket{r,m}=\left[\left(r+m\right)\ldots\left(r+m-p\right)\right.$
×$\left.\times\left(r-m+1\right)\ldots\left(r-m+p\right)\right]^{1/2}\Ket{r,m-p}.$ (69)
and
$\Bra{r,m}\left(\hat{R}^{+}\right)^{p}$ $\left(\hat{R}^{-}\right)^{q}\Ket{r,m}=\left(r+m\right)\ldots\left(r+m-p\right)$
×$\times\left(r-m+1\right)\ldots\left(r-m+p\right)\delta_{pq}.$ (70)
At a given position $z$ the magnetization will be defined by the atoms in the vicinity (see equation 45) and we can approximately replace $\hat{R}^{\pm}$ by $\hat{R}^{\pm}_{k}$ in equation (70). We apply this relation to two different situations.
When $p\ll\left|m\right|\approx r$ we have
$\Braket{r,m}{\left(\hat{R}^{+}_{k}\right)^{p}\left(\hat{R}^{-}_{k}\right)^{q}}{r,m}\simeq\delta_{pq}\,p!\left(2r\right)^{p},$ (71)
which is the situation considered by Gross and Haroche for superradiance with $r=N/2$ [20]. On the other hand, whenever $p\ll\left|m\right|\lesssim r$ equation (70) reduces to
$\Braket{r,m}{\left(\hat{R}^{+}_{k}\right)^{p}\left(\hat{R}^{-}_{k}\right)^{q}}{r,m}\simeq\delta_{pq}\left(r^{2}-m^{2}\right)^{p}.$ (72)
At thermal equilibrium when $\left|m\right|\ll N/2$, calculations are complicated by the very high degeneracy of states. However, we can simplify matters by considering the mean values $\overline{m}$ and $\overline{r\left(r+1\right)}$. Whenever $\hbar\omega_{0}\ll k_{\mathrm{B}}T_{\mathrm{kin}}$, which is the case for the 21 cm line, these are found to be [1]
$\overline{m}$ $\simeq\frac{N}{4}\frac{\hbar\omega_{0}}{k_{\mathrm{B}}T_{\mathrm{kin}}}$ (73)
$\overline{r\left(r+1\right)}$ $\simeq m^{2}+\frac{N}{2},$ (74)
where the latter is valid for any $m$ satisfying our assumption and $k_{\mathrm{B}}$ is the Boltzmann constant. With these relations we can now use equation (72) to find
$\Braket{\overline{r},\overline{m}}{\left(\hat{R}^{+}_{k}\right)^{p}\left(\hat{R}^{-}_{k}\right)^{q}}{\overline{r},\overline{m}}\simeq\delta_{pq}\left(\frac{N}{2}\right)^{p}.$ (75)
Once again following Gross and Haroche [20], we assign the result for an initial measurement on $\hat{R}^{+}_{k}$ to be
$\Braket{\overline{r},\overline{m}}{\hat{R}^{+}_{k}\left(0\right)}{\overline{r},\overline{m}}\equiv\beta e^{i\varphi}$ (76)
with $\beta$ and $\varphi$ random variables. We thus seek probability distributions $P\left(\beta^{2}\right)$ and $Q\left(\varphi\right)$ such that
[…]
with the outcome $P\left(\beta^{2}\right)=\delta\left(\beta^{2}-N/2\right)$, $Q\left(\varphi\right)=1/2\pi$ and $\overline{\beta^{2}}=N/2$.
Further considerations and transformations of the MBE (with $T_{1}=T_{2}\rightarrow\infty$ in equations 39-40) reveal that $\left(\mu n_{-}^{\prime}\right)^{2}+M^{+}M^{-}$ is a conserved quantity over time. This motivates the definition of classical versions for the pseudo-spin operators as
$R^{3}\left(\tau\right)$ $=\frac{\left|N_{-}^{\mathrm{eq}}\right|}{2}\cos\left[\theta\left(\tau\right)\right]$ (78)
$R^{+}_{k}\left(\tau\right)$ $=\frac{\left|N_{-}^{\mathrm{eq}}\right|}{2}\sin\left[\theta\left(\tau\right)\right]e^{i\varphi},$ (79)
which eventually leads to the so-called sine-Gordon equation for superradiance [20, 7].
Setting the initial Bloch angle to $\theta\left(0\right)=\pi-\epsilon$, with $\epsilon\ll 1$, we find from equations (76) and (79) that $\epsilon=2\beta/\left|N_{-}^{\mathrm{eq}}\right|$ and
$\theta_{0}$ $\equiv\sqrt{\overline{\epsilon^{2}}}$
$=\frac{\sqrt{2N}}{\left|N_{-}^{\mathrm{eq}}\right|}$ (80)
for the initial “rising” angle. From equations (45) and (79)-(80), the initial magnetization becomes
$M^{+}\left(z,0\right)=\frac{1}{2}n_{-}^{\mathrm{eq}}\mu\theta_{0}.$ (81)
Appendix C Collision cross-section
We consider a two-atom system initially in the state
[…]
where $\mathbf{P}=\mathbf{p}_{1}+\mathbf{p}_{2}$ and $\Delta\mathbf{p}=\left(\mathbf{p}_{2}-\mathbf{p}_{1}\right)/2$ with $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$ the linear momenta of the first and second atoms, respectively, and the last state |$\Ket{0}$ stands for the vacuum radiation field. We will assume that the two atoms are at resonance, i.e., $\Delta\mathbf{p}=0$ initially. We rewrite the interaction term of the Hamiltonian in equation (2) as follows
$\hat{V}=-i\sum_{q}\frac{1}{2}\hbar\Omega_{q}$ $\sum_{j=1}^{2}\left[\hat{R}_{j}^{+}\hat{T_{q}}\left(\hbar\mathbf{k}_{q}\right)\hat{a}_{q}-\hat{R}_{j}^{-}\hat{T_{q}}^{\dagger}\left(\hbar\mathbf{k}_{q}\right)\hat{a}_{q}^{\dagger}\right]$ (83)
with $\hat{T_{q}}\left(\hbar\mathbf{k}_{q}\right)=e^{i\mathbf{k}_{q}\cdot\hat{\mathbf{r}}}$ the linear momentum translation operator [37].
Using a first-order perturbation expansion, it can be shown that the state of the system after the emission of a first photon is composed of a superposition of states of the form
[…]
that differ only in the radiation mode of the emitted photon. In equation (84) the radiation state |$\Ket{1_{r}}$ implies that one photon occupies mode $r$ and all other modes are empty. After a short time the photon leaves the system and |$\Ket{1_{r}}$ is replaced by |$\Ket{0}$.
We can express this state using the position basis |$\Ket{\mathbf{R},\Delta\mathbf{r}}$, where $\mathbf{R}=\left(\mathbf{r}_{1}+\mathbf{r}_{2}\right)/2$ and $\Delta\mathbf{r}=\mathbf{r}_{2}-\mathbf{r}_{1}$ are conjugate to $\mathbf{P}$ and $\Delta\mathbf{p}$, respectively, to get
|$\Ket{\psi_{+}}$ $\propto\int d^{3}R\,e^{i\left(\mathbf{P}-\hbar\mathbf{k}_{r}\right)\cdot\mathbf{R}}\,\Ket{\mathbf{R}}\int d^{3}\Delta r\,\Ket{\Delta\mathbf{r}}$
×$\times\frac{1}{\sqrt{2}}\left(e^{i\hbar\mathbf{k}_{r}\cdot\Delta\mathbf{r}/2}\Ket{ab}+e^{-i\hbar\mathbf{k}_{r}\cdot\Delta\mathbf{r}/2}\Ket{ba}\right)\Ket{0}.$ (85)
We recognize on the last line the internal state |$\Ket{1,0}_{\theta}$ of equation (8) when $\mathbf{k}_{r}$ makes an angle $\theta$ with $\Delta\mathbf{r}$. The state orthogonal to |$\Ket{\psi_{+}}$, and related to |$\Ket{0,0}_{\theta}$ in equation (10), is
|$\Ket{\psi_{-}}$ $\propto\int d^{3}R\,e^{i\left(\mathbf{P}-\hbar\mathbf{k}_{r}\right)\cdot\mathbf{R}}\,\Ket{\mathbf{R}}\int d^{3}\Delta r\,\Ket{\Delta\mathbf{r}}$
×$\times\frac{1}{\sqrt{2}}\left(e^{i\hbar\mathbf{k}_{r}\cdot\Delta\mathbf{r}/2}\Ket{ab}-e^{-i\hbar\mathbf{k}_{r}\cdot\Delta\mathbf{r}/2}\Ket{ba}\right)\Ket{0}.$ (86)
As discussed in Sec. IV.2, after the transient regime, when the gas containing a large number of such systems has become dark, every two-atom system is in the anti-symmetric dark state given in equation (86). Let us then consider two such systems $A$ and $B$ entering into a collision. We choose a reference frame at rest with the center-of-mass for the two systems such that $\mathbf{R}_{A}=-\mathbf{R}_{B}=-\mathbf{R}_{0}$. For the 21 cm line of the hydrogen atom we can assume that $\left|\mathbf{P}_{A}\right|,\,\left|\mathbf{P}_{B}\right|\gg\hbar k_{r}$ irrespective of the radiation mode. To lighten the notation we omit the spatial states |$\Ket{\mathbf{R}_{A}}$, |$\Ket{\Delta\mathbf{r}_{A}}$, etc., and the vacuum |$\Ket{0}$ state to focus on the internal states, such that we can write for the colliding system
|$\Ket{\Psi}$ =|$=\Ket{\psi_{-}}_{A}\Ket{\psi_{-}}_{B}$
$\propto e^{i\left(\mathbf{P}_{B}-\mathbf{P}_{A}\right)\cdot\mathbf{R}_{0}/\hbar}\left(\Ket{ab}_{A}-\Ket{ba}_{A}\right)\left(\Ket{ab}_{B}-\Ket{ba}_{B}\right)$
[…]
If two atoms enter into an elastic collision, say the first atoms of both pairs, then under the Born approximation each term will bring a scattering amplitude [38]
$f\left(\mathbf{k}_{\mathrm{in}},\mathbf{k}_{\mathrm{out}}\right)=-\frac{m_{\mathrm{r}}}{2\pi\hbar^{2}}\int d^{3}r^{\prime}e^{-i\mathbf{k}_{\mathrm{out}}\cdot\mathbf{r}^{\prime}}V\left(\mathbf{r}^{\prime}\right)e^{i\mathbf{k}_{\mathrm{in}}\cdot\mathbf{r}^{\prime}}$ (88)
with $\mathbf{k}_{\mathrm{in}}=\left(\mathbf{P}_{B}-\mathbf{P}_{A}\right)/\hbar$, $\mathbf{k}_{\mathrm{out}}$ the scattered wave vector and $m_{\mathrm{r}}$ the reduced mass. It therefore follows from equation (87) that the collision cross-section $\sigma\left(\mathbf{k}_{\mathrm{in}},\mathbf{k}_{\mathrm{out}}\right)=0$ since it consists of the square of the sum of the four corresponding scattering amplitudes.
Although such calculations need to be generalized or formalized, we find that two darkened entangled systems are likely to pass through each other, unimpeded, during a collision. They would not feel the presence of each other even though individual atoms may actually collide. This is due to the destructive interference resulting from the entanglement within each of the systems.
Finally, it is interesting to note that in cases where the states are totally symmetric the cross-section scales as $N^{2}\sigma_{0}$, where $N$ is the number of atoms and $\sigma_{0}$ the one-atom cross-section.